# An Explicit Wavefunction of the Interacting Non-Hermitian Spin-1/2 1D System

### v1

@@ Abstract
in which $\beta=1$ corresponds to the realistic situation described by Eq. (20). Consider the zero-temperature limit $\beta\rightarrow\infty$ such that the system freeze into the minimal energy configuration of $\mathcal{H}$. As a simple example, we examine the case of 4 particles shown in SM. C Fig. 4. The energy minima at small and large values of $\eta_{r}$ are calculated, which correspond to frozen configurations shown in Fig. 2 ($c$) and ($d$) respectively. The switch of minima occurs at $\eta_{r}=3.84$, roughly matching the transition point shown in Fig. 2 ($f$).
As for experimental realizations, consider the following 1D lattice Hamiltonian,
[…]
$+$ $i\gamma\sum_{n,\alpha}(-1)^{n}c_{n,\alpha}^{\dagger}c_{n,\alpha},$
where $t_{1}$, $t_{2}$ and $\gamma$ are real, and $|t_{1}|>|t_{2}|$; the $t_{2}$-term stands for the Hermitian SOC; the $\gamma$-term represents the on-site loss, which can be realized in cold atomic systems [18]. Although Eq. (23) is different from our non-Hermitian SOC, it exhibits the similar spin-dependent NHSE [41], i.e., spin-up and down particles localize at opposite boundaries upon the OBC. It would be interesting to further investigate the behavior of such a system once interactions are turned on. Due to the appearance of the next-nearest-neighbor hopping, such a system is no longer integrable. Nevertheless, our results via BA provide a good starting point for further exploring the exotic physics on the interplay between strong interaction and non-Hermitian physics.
Discussion and conclusion.— We present a BA solution to a 1D interacting spin-$\frac{1}{2}$ non-Hermitian system breaking the inversion symmetry. The interplay between non-Hermitian SOC and the repulsive interaction results in a novel many-body resonance state. The complicated BA wavefunction is simplified in the dilute limit, which is cast into a product between the Slater determinants and a Jastrow factor. Its amplitude square is mapped into a thermodynamic distribution, exhibiting the competition between Pauli’s exclusion among fermions of the same component and resonances between fermions of different components. The former brings a repulsion and the latter generates an unidirectional attraction. This competition leads to a transition from a uniform configuration with weak pairing tendency to a phase separation.
@@ Acknowledgements.
We are grateful to K. Yang, W. Yang and C. H. Ke for valuable discussions. Z. Y. is supported by the National Key Research and Development Program of China (Grant No. 2023YFA1407500), the National Natural Science Foundation of China (12322405, 12104450, 12047503), and the Fundamental Research Funds for the Central Universities (20720230011). C.W. is supported by the National Natural Science Foundation of China under the Grants No. 12234016 and No. 12174317. This work has been supported by the New Cornerstone Science Foundation.
Appendix A Exact solution of two-body problem
@@ Appendix B Solution to BA equation
×$\times A_{\sigma}(Q,P)\ \exp\left(i\sum_{l}\overline{k}_{P_{l}}x_{Q_{l}}\right),$
where $Q$ and $P$ represents permutations of the coordinates $\{x_{l}\}$ and momenta $\{\overline{k}_{l}\}$ respectively. $\sum_{Q,P}$ denotes a sum over all permutations. For eigenstates with $N_{\downarrow}$ and $N_{\uparrow}=N-N_{\downarrow}$ spin-down and spin-up particles, the corresponding BA equations are [39]
@@
×$\times A_{\sigma}(Q,P)\ \exp\left(i\sum_{l}{k}_{P_{l},\sigma_{l}}x_{Q_{l}}\right).$
In this expression, $k_{P_{l},\sigma_{l}}=\overline{k}_{P_{l}}-im\alpha\sigma_{l}$ with $\sigma_{l}=\uparrow,\downarrow$. Without loss of generality, we assume $\sigma_{i}=\downarrow$, $\sigma_{j}=\uparrow$ for any $1\leq i\leq N_{\downarrow}$, $N_{\downarrow}+1\leq j\leq N$. For a given permutation $\{P_{l}\}$ and $\{Q_{l}\}$, $\{k_{P_{l},\sigma_{l}}\}$ defines a scattering channel, where $1\leq P_{l}\leq N$. These scattering channels can be divided into two classes:
[…]
We will now prove that class (a) can be discarded in the dilute limit. For a channel in class (a), if the momentum of a spin-down particle satisfies $\mbox{Im}\ k_{P_{i},\sigma_{i}}=2m\alpha-N_{\uparrow}\eta_{r}/L$, there must exist a corresponding spin-up particle with momentum $\mbox{Im}\ k_{P_{j},\sigma_{j}}=-(2m\alpha-N_{\downarrow}\eta_{r}/L)$. Class (a) can have many such up-down pairs, each pair contributing an exponential factor $e^{2(x_{\uparrow}-x_{\downarrow})/\lambda_{s}}$. Depending on the coordinate permutation ($-L<x_{\uparrow}-x_{\downarrow}<0$ or $0<x_{\uparrow}-x_{\downarrow}<L$), this factor is either divergent or suppressed in the case of $L\gg\lambda_{s}$. If it diverges, the PBC can not be satisfied unless the corresponding scattering amplitude scales as $e^{-2L/\lambda_{s}}$. Therefore, regardless of whether the plane waves diverge or not, class (a) will always be suppressed by a factor $e^{-2x/\lambda_{s}}$ with $0<x<L$. An exception occurs when $x<\lambda_{s}$, which means a pair of spin-up and spin-down particles get very close to each other. However, in the dilute limit, the average distance between particles $L/N$ is much larger than $\lambda_{s}$, rendering the exceptional region negligible. Furthermore, scattering channels of class (a) will not accumulate to a finite value as $N\rightarrow\infty$. Although the number of these channels increase as $N!$, they are not coherent and will cancel out each other. To justify this argument, we numerically calculate the normalization of the approximated wavefunction and the exact wavefunction in the 4-particle case with 2 up-down pairs. They differ by a ratio of the magnitude $N\lambda_{s}/L$, which is roughly the normalization of a single channel in class (a).
@@
Figure 4: In the first configuration, four particles form two up-down pairs with distance $L/2$. The size of the pair is negligible; In the phase separation, particles with identical spins form two adjacent clusters with size $a=L\tan^{-1}(\frac{\pi}{2}\eta_{r})/\pi$.
@@
As we sum over all momentum permutations in the wavefunction Eq. (33), only those channels of class (b) should be included. For any fixed coordinate permutation, different channels in class (b) can transform to each other by reflection between particles with identical spins, and the corresponding phase shift is $-1$ due to Pauli’s exclusion principle. Consequently, these plane waves with momentum distribution Eq. (34) organize into a product of two Slater determinants for two spin components, simplifying the wavefunction to:
@@
×$\times\sum_{Q}\theta\left(x_{Q_{1}}>x_{Q_{2}}>\cdots>x_{Q_{N}}\right)A_{\sigma}(Q,Q)$
@@
The function $A_{\sigma}(x_{1},x_{2},\cdots x_{N})$ is a constant for a given coordinate permutation. Its value changes only when the coordinates of an up-down pair switch. When all the spin-up particles are on the left side of spin-down particles, its value is defined as 1. According to the constant phase shift Eq. (37), every time a spin-up particle cross a spin-down particle from the left to the right, the value of $A_{\sigma}(x_{1},x_{2},\cdots x_{N})$ will multiply a factor $e^{-\eta_{r}}$. Therefore, we can use the following algorithm to determine the output of the function:
@@
Appendix C Phase transition in the four-particle case
@@
$\rho_{\beta}=e^{-\beta\mathcal{H}}.$ (40)
@@
In the zero-temperature limit $\beta\rightarrow\infty$, the system freezes into the state with minimal energy. For the four-particle case, $\mathcal{H}$ has two local minima, whose configurations are shown in Fig. 4. The energies of the uniform configuration $E_{1}$ and the phase separation $E_{2}$ are given by:
$\left\{\begin{array}[]{l}E_{1}=2\eta_{r},\vspace{1ex}\\ E_{2}=\frac{8\eta_{r}a}{L}-4\ln(\sin\frac{\pi a}{L}).\vspace{1ex}\end{array}\right.$ (43)
@@
As a function of the cluster size $a$, $E_{2}$ is minimized when $a=L\tan^{-1}(\frac{\pi}{2}\eta_{r})/\pi$. The transition point is determined by setting $E_{1}=E_{2}$, which gives $\eta_{r}=3.84$. In the case of $\eta_{r}<3.84$ or $\eta_{r}>3.84$, $E_{1}$ or $E_{2}$ is the smaller one.

### v2

@@ Abstract
in which $\beta=1$ corresponds to the situation described by Eq. (18). Consider the zero-temperature limit $\beta\rightarrow\infty$ such that the system freeze into the minimal energy configuration of $\mathcal{H}$. As an example, we examine the case of 4 particles shown in SM. III. The energy minima at small and large values of $\eta_{r}$ are calculated, which correspond to frozen configurations shown in Fig. 3 ($a$) and ($b$) respectively. The switch of minima occurs at $\eta_{r}=3.84$, roughly matching the transition point shown in Fig. 3.
Discussion and conclusion.— We present a BA solution to a 1D interacting spin-$\frac{1}{2}$ non-Hermitian system breaking the inversion symmetry. The interplay between non-Hermitian SOC and the repulsive interaction results in a novel many-body resonance state. The complicated BA wavefunction is simplified in the dilute limit, which is cast into the Slater-Jastrow form. Its amplitude square is mapped into a thermodynamic distribution, exhibiting the competition between Pauli’s exclusion among fermions of the same component and resonances between fermions of different components. The former brings a repulsion and the latter generates an unidirectional attraction. This competition leads to a transition from a uniform configuration with weak pairing tendency to a phase separation, showing that the many-body resonance is enhanced by the repulsive interaction.
@@ Acknowledgements.
We are grateful to K. Yang, W. Yang, J. D. Wu and C. H. Ke for valuable discussions. Z. Y. is supported by the National Key Research and Development Program of China (Grant No. 2023YFA1407500), the National Natural Science Foundation of China (12322405, 12104450, 12047503), and the Fundamental Research Funds for the Central Universities (20720230011). C.W. is supported by the National Natural Science Foundation of China under the Grants No. 12234016 and No. 12174317. This work has been supported by the New Cornerstone Science Foundation.
Supplemental Material
II I. Exact solution of two-body problem
@@
$\times A_{\sigma}(Q,P)\ \exp\left(i\sum_{l}\overline{k}_{P_{l}}x_{Q_{l}}\right),$
where $Q$ and $P$ represents permutations of the coordinates $\{x_{l}\}$ and momenta $\{\overline{k}_{l}\}$ respectively. $\sum_{Q,P}$ denotes a sum over all permutations. For eigenstates with $N_{\downarrow}$ and $N_{\uparrow}=N-N_{\downarrow}$ spin-down and spin-up particles, the corresponding BA equations are
@@
$\times A_{\sigma}(Q,P)\ \exp\left(i\sum_{l}{k}_{P_{l},\sigma_{l}}x_{Q_{l}}\right).$
In this expression, $k_{P_{l},\sigma_{l}}=\overline{k}_{P_{l}}-im\alpha\sigma_{l}$ with $\sigma_{l}=\pm 1$ representing the spin $z$-component. Without loss of generality, we assume $\sigma_{i}=-1$, $\sigma_{j}=1$ for any $1\leq i\leq N_{\downarrow}$, $N_{\downarrow}+1\leq j\leq N$. For a given permutation $\{P_{l}\}$ and $\{Q_{l}\}$, $\{k_{P_{l},\sigma_{l}}\}$ defines a scattering channel, where $1\leq P_{l}\leq N$. These scattering channels can be divided into two classes:
$\text{Class (a):}\quad\exists\ l,\ \mbox{Im}\ k_{P_{l},\sigma_{l}}$ $=\pm(2m\alpha-N_{\downarrow,\uparrow}\eta_{r}/L),$
$\text{Class (b):}\quad\forall\ l,\ \mbox{Im}\ k_{P_{l},\sigma_{l}}$ $=\pm N_{\uparrow,\downarrow}\eta_{r}/L.$
We will now prove that class (a) can be discarded in the dilute limit. For a channel in class (a), if the momentum of a spin-down particle satisfies $\mbox{Im}\ k_{P_{i},\sigma_{i}}=2m\alpha-N_{\uparrow}\eta_{r}/L$, there must exist a corresponding spin-up particle with momentum $\mbox{Im}\ k_{P_{j},\sigma_{j}}=-(2m\alpha-N_{\downarrow}\eta_{r}/L)$. Class (a) can have many such up-down pairs, each pair contributing an exponential factor $e^{2(x_{\uparrow}-x_{\downarrow})/\lambda_{s}}$. Depending on the coordinate permutation ($L<x_{\uparrow}-x_{\downarrow}<0$ or $0<x_{\uparrow}-x_{\downarrow}<L$), this factor is either divergent or suppressed in the case of $L\gg\lambda_{s}$. If it diverges, the PBC can not be satisfied unless the corresponding scattering amplitude scales as $e^{-2L/\lambda_{s}}$. Therefore, regardless of whether the plane waves diverge or not, class (a) will always be suppressed by a factor $e^{-2x/\lambda_{s}}$ with $0<x<L$. An exception occurs when $x<\lambda_{s}$, which means a pair of spin-up and spin-down particles get very close to each other. However, in the dilute limit, the average distance between particles $L/N$ is much larger than $\lambda_{s}$, rendering the exceptional region negligible. Furthermore, scattering channels of class (a) will not accumulate to a finite value as $N\rightarrow\infty$. Although the number of these channels increase as $N!$, they are not coherent and will cancel out each other. To justify this argument, we numerically calculate the normalization of the approximated wavefunction and the exact wavefunction in the 4-particle case with 2 up-down pairs. They differ by a ratio of the magnitude $N\lambda_{s}/L$, which is roughly the normalization of a single channel in class (a).
@@
As we sum over all momentum permutations in the wavefunction Eq. (30), only those channels of class (b) should be included. For any fixed coordinate permutation, different channels in class (b) can transform to each other by reflection between particles with identical spins, and the corresponding phase shift is 1-1 due to Pauli’s exclusion principle. Consequently, these plane waves with momentum distribution Eq. (31) organize into a product of two Slater determinants for two spin components, simplifying the wavefunction to:
@@
$\times\sum_{Q}\theta\left(x_{Q_{1}}>x_{Q_{2}}>\cdots>x_{Q_{N}}\right)A_{\sigma}(Q,Q)$
@@
The function $A_{\sigma}(x_{1},x_{2},\cdots x_{N})$ is a constant for a given coordinate permutation. Its value changes only when the coordinates of an up-down pair switch. When all the spin-up particles are on the left side of spin-down particles, its value is defined as 1. According to the constant phase shift Eq. (34), every time a spin-up particle cross a spin-down particle from the left to the right, the value of $A_{\sigma}(x_{1},x_{2},\cdots x_{N})$ will multiply a factor $e^{-\eta_{r}}$. Therefore, we can use the following algorithm to determine the output of this function:
@@
Figure 5: ($a$) and ($b$) shows the Pauli repulsion energy $V(r)$ and the resonance energy $W(r)$ respectively. They are periodical functions due to the periodical boundary condition. ($c$) illustrates the configurations with the locally minimal energy in the zero temperature limit. In the first one, four particles form two up-down pairs with distance $L/2$. The size of the pair is negligible; In the second one, particles with identical spins form two adjacent clusters with size $a=L\tan^{-1}(\frac{\pi}{2}\eta_{r})/\pi$.
IV III. Phase transition in the four-particle case
@@
$\rho_{\beta}=e^{-\beta\mathcal{H}},$ (37)
@@
where $\mathcal{H}$ is a sum over the Pauli repulsion energy $V(r)$ between particles of identical spins and the resonance energy $W(r)$ between particles of different spins. The shape of $V(r)$ and $W(r)$ is illustrated in Fig. 5 ($a$) and ($b$) respectively. In the zero-temperature limit $\beta\rightarrow\infty$, the system freezes into the state with minimal energy. For the four-particle case, $\mathcal{H}$ has two local minima, whose configurations are shown in Fig. 5. The energies of the uniform configuration $E_{1}$ and the phase separation $E_{2}$ are given by:
$\left\{\begin{array}[]{l}E_{1}=2\eta_{r},\vskip 4.30554pt\\ E_{2}=\frac{8\eta_{r}a}{L}-4\ln(\sin\frac{\pi a}{L}).\vskip 4.30554pt\end{array}\right$. (40)
@@
As a function of the cluster size $a$, $E_{2}$ is minimized when $a=L\tan^{-1}(\frac{\pi}{2}\eta_{r})/\pi$. The transition point is determined by setting $E_{1}=E_{2}$, which gives $\eta_{r}=3.84$. One finds the energy minimum switch from $E_{1}$ to $E_{2}$ as crossing the transition point, which means the uniform configuration turns to phase sepration, as prediction.
Figure 6: Lattice chain corresponding to the Hamiltonian $\hat{H}_{L}$. Each unit cell of the chain consists of two sublattice sites, denoted by green and orange dots respectively. Dissipation is introduced on the orange sites via coupling to a reservoir, which is modeled by the on-site loss term $i\gamma c_{n}^{\dagger}c_{n}$, while the green sites are free of loss. A spin-dependent flux $\phi=\sigma\pi/2$ threads each triangular plaquette formed by the green and orange sites, contributing to a phase shift to the next nearest neighbour hopping. $\sigma=\pm 1$ represents the spin $z$-component.
V IV. Experimental realization
Figure 7: ($a$) shows all the single-body eigenstates of the lattice Hamiltonian $\hat{H}_{L}$ upon the OBC. $n$ and $\psi_{n}$ represent the index of the lattice site and the value of the eigenstate $\psi$ at this site, respectively. As illustrated, spin-up and down particles localize at different boundaries. Here $t_{1}=1$, $t_{2}=0.1$ and $\gamma=1$. ($b$) and ($c$) shows the moving patterns of spin-up particles ($\sigma=1$) on the lattice defined by $\hat{H}_{l}$. ($b$) Right and left moving modes tend to take different paths due to the magnetic flux. ($c$) The loss on the orange site does not affect the main path of the right moving mode, while blocks that of the left moving mode.
As for experimental realizations, consider the following 1D lattice Hamiltonian
$\begin{aligned} \hat{H}_{L}&=-\sum_{n,\sigma}\left(t_{1}c_{n,\sigma}^{\dagger}c_{n+1,\sigma}-it_{2}(-1)^{n}c_{n,\sigma}^{\dagger}\sigma^{z}c_{n+2,\sigma}+\text{h.c.}\right)\\ &-i\gamma\sum_{n\text{ odd},\ \sigma}(-1)^{n}c_{n,\sigma}^{\dagger}c_{n,\sigma}.\end{aligned}$
Here $n$ and $\alpha$ of the creartion operator $c_{n,\alpha}^{\dagger}$ are the indices for site and spin, respectively. $t_{1}$, $t_{2}$, $\gamma$ are all real, with $|t_{1}|>|t_{2}|$. $t_{2}$-term stands for the Hermitian SOC; $\gamma$-term represents the loss on the odd index sites, which can be realized in the cold atomic systems [21]. Notice that $H_{L}$ is free of gain, in which case the dynamics described the Lindbladian can be fully determined by its corresponding non-Hermitian Hamiltonian [3]. In Fig. 6 we illustrate the corresponding lattice chain with green and orange dots denoting different sublattice sites.
The lattice Hamiltonian $\hat{H}_{L}$ and the continuum one $\hat{H}$ are equivalent in the sense that $\hat{H}_{L}$ exhibits an effective imaginary spin-orbit coupling (SOC). One can check this in the momentum representation of $\hat{H}_{L}$, given by
$\mathcal{H}_{L}(k)=\left(\begin{array}[]{cc}2t_{2}\sigma^{z}\sin k&t_{1}\left(1+e^{-ik}\right)\\ t_{1}\left(1+e^{ik}\right)&-2t_{2}\sigma^{z}\sin k-i\gamma\end{array}\right),$ (41)
Here $\mathcal{H}_{L}(k)$ is written in the sublattice space. The eigen-value of $\mathcal{H}_{L}(k)$ is
$E_{k}=-i\frac{\gamma}{2}\pm\sqrt{\left(2t_{1}\cos\frac{k}{2}\right)^{2}+\left(2t_{2}\sin k+i\frac{\gamma}{2}\sigma\right)^{2}},$
where $\sigma=\pm 1$ represents the spin $z$-component. In the long-wavelength limit, $E_{k}$ can be expanded in powers of $k$ up to second order:
$E_{k}=-i\frac{\gamma}{2}\pm\left(\epsilon_{0}+\frac{k^{2}+2im\alpha\sigma k}{2m}\right),$ (42)
with
$m=\frac{\epsilon_{0}}{t_{1}^{2}}\left[\left(\frac{4t_{2}}{\epsilon_{0}}\right)^{2}-1\right]^{-1},\quad\alpha=\frac{\gamma t_{2}}{\epsilon_{0}},\quad\epsilon_{0}=\sqrt{4t_{1}^{2}-\frac{\gamma^{2}}{4}}.$
In this limit, the lattice Hamiltonian $\hat{H}_{L}$ reduces to the continuous Hamiltonian $\hat{H}$, where the term $2im\alpha\sigma k$ in $\hat{H}_{L}$ captures the non-Hermitian SOC. Although this non-Hermitian term contributes an unbounded imaginary part to $E_{k}$ at the level of perturbative expansion, the exact non-perturbative expression of $E_{k}$ indicates that $\text{Im}(E_{k})<0$ for all $k$, implying that the system intrinsically has no gain. Furthermore, exact diagonalization of $\hat{H}_{L}$ in the case of open boundary condition shows the spin-dependent NHSE, i.e., eigenstates of spin-up and down particles localize at different boundaries. This is shown in Fig 7 ($a$).
The spin-dependent NHSE in this lattice model can be explained as following. We start from a tight bonding chain with a single triangular cell in the middle, and take $\sigma=1$. As illustrated in Fig 7. ($b$), when the orange site is free of loss, the Hamiltonian of the system is
$\hat{H_{l}}=\hat{H}_{0}+\hat{H}_{c}\ ,\quad\hat{H}_{0}=-\sum_{n}it_{2}(c_{n}^{\dagger}c_{n+2}-c_{n+2}^{\dagger}c_{n})\ ,$
$\hat{H}_{c}=-t_{1}(d^{\dagger}c_{m}+d^{\dagger}c_{m+1}+\text{h.c.})\ ,$
where $\hat{H}_{0}$ represents the Hamiltonian of the tight bonding chain composed by the green sites, and $\hat{H}_{c}$ describes the coupling between the chain and the orange site, with $d^{\dagger}$ denoting the creation operator on that site. Under this setup, due to the interference of the wavefunction, right moving plane waves will have lower occupation on the orange site than that of the left moving ones. For example, a right moving plane wave with momentum $k=\pi$ does not pass through the orange site. It moves to the right because $v_{k}=dE_{0}/dk$ is positive at $k=\pi$, where $E_{0}$ is the eigen-value of $\hat{H}_{0}$; and its occupation vanishes on the orange site because the hoppings from $m$th and $(m+1)$th sites cancel with each other. On the contrary, coherent interference happens on the orange site for left moving plane waves, for instance with $k=0$. Therefore, when the orange site is subjected to loss, right moving waves can avoid it while left moving waves can not, such that the life time of the former will be longer than that of the latter. Consequently, spin-up particles will accumulate on the right boundary upon the OBC. This is illustrated in Fig. 7 ($c$). The reverse skin effect happens for spin-down particles, which can be understood by performing a parity transformation to $\hat{H}_{l}$. The lattice model given by $\hat{H}_{L}$ is obtained by repeating the triangular cell described above. In this case, all the long-live modes mainly distribute on the green chain to avoid the loss on the orange chain, where spin-up and down particles preferring right and left hopping respectively. This induces an effective non-Hermitian spin-dependent hopping, giving rise to the spin-dependent NHSE.
With the imaginary SOC, it can be expected that $\hat{H}_{L}$ will also exhibits many-body resonance effect if the repulsive interaction is turned on. Due to the appearance of the next-nearest-neighbor hopping in $\hat{H}_{L}$, such a system is no longer integrable. Nevertheless, our results via BA provide a good starting point for further exploring the exotic physics on the interplay between strong interaction and non-Hermitian physics.
The resonance effect can be experimental detected by studying the dynamics of particles, which is similar to the observation of single-body NHSE. In the single-body case, a spin-up particle initially placed in the middle of the lattice chain will propagate to the right side due to the NHSE. If a spin-down particle is added on its right side, the pair will form a stable state in the bulk instead of propagating to the opposite sides, which can be seen as an evidence for the resonance effect. By placing many spin-up and down particles on the chain, the two groups of particles will mix when the repulsive interaction is weak and become spatially separated if the interaction strength exceeds a critical value, which is the signal for the phase transition.